OpenStudy (anonymous):

Find the inverse function of f(x)=1/x+5 f-1(x)=__________________? help

OpenStudy (anonymous):

f^-1(x) = (1/x) - 5

OpenStudy (anonymous):

@royrdz look at other answer it was wrong

OpenStudy (anonymous):

be careful here ok?

OpenStudy (anonymous):

it is not \[\frac{1}{x}-5\] it is \[\frac{1}{x-5}\]

OpenStudy (anonymous):

First of all you must know that not all functions have inverses. Therefore to verify you must us the following statement: if f(a)=f(b) then a=b Now just plug in the statement and you have that \[1/a +5 = 1/b +5\] which by simple arithmetic you have that a=b which states that the function has an inverse. Now to find the inverse you just substitute x by f−1(x) and f(x) by x which in this case is the following: \[x = 1/f^{-1}(x) + 5\] and clear for the inverse function: \[x - 5 = 1/f ^{-1}(x)\] ; \[1/(x-5) = f ^{-1}(x)\] which is the result you needed. Hope it helps form the beautiful island of Puerto Rico. Peace, Love and Happiness

OpenStudy (anonymous):

\[y=\frac{1}{x}+5\] switch x and y get \[x=\frac{1}{y}-5\] solve for y get \[x+5=\frac{1}{y}\] \[y=\frac{1}{x+5}\] so \[f^{-1}=\frac{1}{x-5}\]

OpenStudy (anonymous):

what jobh said

OpenStudy (anonymous):

\[f-1 ( x ) = (1\div x) -5\]

OpenStudy (anonymous):

\[f ^{-1}(y)=(1-5y)/y, i.e. f ^{-1}(y)=(1/y)-5\]